OpenStudy (anonymous):

Can someone try and help me with this Find the sum twelve and three-eighths plus two and one-fifth. Be sure to reduce to simplest form. fourteen and four-thirteenths fourteen and twenty-three fortieths twelve and twenty-three fortieths fourteen and four-fifths

jigglypuff314 (jigglypuff314):

\[12\frac{ 3 }{ 8 } + 2\frac{ 1 }{ 5 } = 12 + 2 + \frac{ 3 }{ 8 } + \frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

14 4/20

OpenStudy (anonymous):

4/13

OpenStudy (anonymous):

right @jigglypuff314 ?

jigglypuff314 (jigglypuff314):

mmm try again? 3 1 15 8 15 + 8 -- +-- = -- + -- = -------- = ? 8 5 40 40 40

OpenStudy (anonymous):

23/40 but are you able to simply it

jigglypuff314 (jigglypuff314):

23/40 is the simplest you can get it :)

OpenStudy (anonymous):

thank you @jigglypuff314

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

cna you help me with one more plz

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

can*

OpenStudy (anonymous):

Find the difference ten and two-ninths minus seven and two-thirds. Be sure to reduce to simplest form. four and five-ninths three and zero-sixths two and five-ninths three and four-ninths

jigglypuff314 (jigglypuff314):

\[10\frac{ 2 }{ 9 } - 7\frac{ 2 }{ 3 } = (10 - 7) + (\frac{ 2 }{ 9 } - \frac{ 2 }{ 3 })\]

OpenStudy (anonymous):

well 10-7 = 3 and It is 0/3

jigglypuff314 (jigglypuff314):

... mmm... well 2 2 2 6 -4 -- - -- = -- - -- = -- 9 3 9 9 9 errr then 3 = 2 + 9/9 right? so 3 (-4/9) = 2 + 9/9 - 4/9

OpenStudy (anonymous):

5/9 right

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

is it 5 5/9

OpenStudy (anonymous):

2

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

Find the sum four-fifteenths plus one-fifth. Be sure to reduce to simplest form. this is 1/3 right

jigglypuff314 (jigglypuff314):

not quite :) try again? 4 1 4 3 4 + 3 -- + -- = -- + -- = ----- = ? 15 5 15 15 15

OpenStudy (anonymous):

7/15

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Find the sum negative fifty one hundredths plus negative twenty-five one hundredths. Be sure to reduce to simplest form. did we do it

OpenStudy (anonymous):

this one?

jigglypuff314 (jigglypuff314):

mmm we haven't done this one before, if that's what you're asking :)

OpenStudy (anonymous):

well can you help me with it?

jigglypuff314 (jigglypuff314):

-51 -25 -51 - 25 ---- + ---- = --------- = ? 100 100 100

OpenStudy (anonymous):

It is -75/100

jigglypuff314 (jigglypuff314):

which can be simplified because -75 and 100 can both be divided by 25

OpenStudy (anonymous):

right